OpenStudy (mieshapitts):

im in highschool and struggling to keep up with being the youngest in my class because im a grade lower than the rest. I just need help on how to graph a parabola with only information about what the vertex and the y intercept/ coordinates are. thanks

OpenStudy (lowkey.s):

okay aany ideas?

OpenStudy (mieshapitts):

like an example?

OpenStudy (lowkey.s):

http://www.csun.edu/~ayk38384/notes/mod11/Parabolas.html this might help

OpenStudy (mieshapitts):

thanks so much

OpenStudy (tkhunny):

If you have ONLY the Vertex and y-intercept, you have no unique solution. The best you can do it two cases. If we ASSUME that the axis of symmetry is parallel to the y-axis, we have the following: 4p(y-k) = (x-h)^2 If the vertex is (1,2), we have 4p(y-2) = (x-1)^2 If the y-intercept is (0,5), we have 4p(5-2) = (0-1)^2 Now, try it assuming the axis of symmetry is parallel to the x-axis.